OpenStudy (anonymous):

Quadratic Formula? 5x^2 + 2x + 2 =

OpenStudy (ash2326):

We have \[5x^2+2x+2=0\] Standard quadratic equations \[ax^2+bx+c=0\] The roots are given by quadratic formula \[x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\] @bigcoxy Can you tell me the value of a, b and c for our equation?

OpenStudy (anonymous):

a =5 b=2 c=2

OpenStudy (ash2326):

Great:) so \[x=\frac{-2\pm \sqrt{(2)^2-4 \times 5 \times 2}}{2 \times 5}\] we get \[x=\frac{-2 \pm \sqrt{-36}}{10}\] we know \[\sqrt{-1}=i\] so \[x=\frac{-2 \pm 6i}{10}\] we get finally \[x=\frac{-1\pm 3i}{5}\]